# Not All Language Model Features Are Linear

### v1

@@ 4 Sparse Autoencoders Find Multi-Dimensional Features
Sparse autoencoders (SAEs) deconstruct model hidden states into sparse vector sums from an over-complete basis [6, 9]. For hidden states $X_{i,l}$, a one-layer SAE of size $m$ with sparsity penalty $\lambda$ minimizes the following dictionary learning loss [6, 9]:
@@
We propose that SAEs can discover irreducible multi-dimensional features by locating point sets with a low $\epsilon$-mixture index. For instance, if $X_{i,l}$ contains an irreducible two-dimensional feature ${\mathbf{f}}$ (see 2), and ${\mathbf{D}}$ includes just two dictionary elements spanning ${\mathbf{f}}$, both elements must have nonzero activations post-ReLU to perfectly reconstruct ${\mathbf{f}}$ (otherwise ${\mathbf{f}}$ is a mixture). Thus the Jaccard similarity of the sets of tokens that these two dictionary elements fire on is likely to be high. On the other hand, if ${\mathbf{D}}$ contains more than two dictionary elements that span ${\mathbf{f}}$, then the Jaccard similarity may be lower. However, since there are now many dictionary elements with a high projection in the two dimensional feature space, the cosine similarity of the dictionary elements is likely to be high.
Thus for a two dimensional irreducible feature ${\mathbf{f}}$, we expect there to be groups of dictionary elements with either high cosine or Jaccard similarity corresponding to ${\mathbf{f}}$. We expect this observation to be true for higher dimensional irreducible features as well. Note that some clusters may correspond to separable features ${\mathbf{f}}$, as this technique only finds features that are not mixtures. This suggests a natural approach to using sparse autoencoders to search for irreducible multi-dimensional features:
1. Cluster dictionary elements by their pairwise cosine similarity or Jaccard similarity.
@@
3. Examine the resulting reconstructed activation vectors for irreducible multi-dimensional features, especially ensuring that the reconstruction is not separable. This step can be done manually by visually inspecting the PCA projections for known irreducible multi-dimensional structures (e.g. circles, see Fig. 2) or automatically by passing the PCA projections to the tests for Definition 3.
@@
This method succeeds on toy datasets of synthetic irreducible multi-dimensional features; see Appendix D. [^1] Code for reproducing experiments: https://github.com/JoshEngels/MultiDimensionalFeatures We apply this method to language models using GPT-2 [41] SAEs trained by Bloom for every layer [5] and Mistral 7B [21] SAEs that we train on layers 8, 16, and 24 (training details in Appendix E). Clustering details are in Appendix F, including comments on scalability.
@@ 5 Circular Representations in Large Language Models
In this section, we seek tasks in which models use the multi-dimensional features we discovered in Section 4, thereby providing evidence that these representations are indeed the fundamental unit of computation for some problems. Inspired by prior work studying circular representations in modular arithmetic [27], we define two prompts that represent “natural” modular arithmetic tasks:
@@
For Weekdays, we range over the $7$ days of the week and durations between $1$ and $7$ days to get $49$ prompts. For Months, we range over the $12$ months of the year and durations between $1$ and $12$ months to get $144$ prompts. Mistral 7B and Llama 3 8B [2] achieve reasonable performance on the Weekdays task and excellent performance on the Months task (measured by comparing the highest logit valid token against the ground truth answer), as summarized in Table 1. Interestingly, although these problems are equivalent to modular arithmetic problems $\alpha+\beta\equiv\;?\ (\mathrm{mod}\ m)$ for $m=7,12$, both models get trivial accuracy on plain modular addition prompts, e.g. “5+3(mod $5+3\ (\mathrm{mod}\ 7)\equiv$”. Finally, although GPT-2 has circular representations, it gets trivial accuracy on Weekdays and Months.
@@ 5.1 Intervening on Circular Day and Month Representations
We first confirm that Llama 3 8B and Mistral 7B have circular representations of $\alpha$ by examining the PCA of hidden states across prompts at various layers on the $\alpha$ token. We plot two of these in Fig. 3 and show all layers in Fig. 14 in the appendix. These plots show circular representations as the highest varying two components in the model’s representation of $\alpha$ at many layers.
@@
We now experiment with intervening on these circular representations. We base our experiments on the common interpretability technique of activation patching, which replaces activations from a “dirty” run of the model with the corresponding activations from a “clean” run [48]. Activation patching empirically tests whether a specific model component, position, and/or representation has a causal influence on the model’s output. We employ a custom subspace patching method to allow testing for whether a specific circular subspace of a hidden state is sufficient to causally explain model output. Specifically, our patching technique relies on the following steps (visualized in Fig. 5):
@@
2. Intervene on the subspace: Say our initial prompt had $\alpha=\alpha_{j}$ and we are intervening with $\alpha=\alpha_{j^{\prime}}$. In this step, we replace the model’s projection on the subspace ${\mathbf{P}}\cdot{\mathbf{W}}_{i,l}$, which will be close to $\texttt{circle}(\alpha_{j}$), with the “clean” point $\texttt{circle}(\alpha_{j^{\prime}})$. Note that we do not use the hidden state $\mathbf{x}^{j^{\prime}}_{i,l}$ from the “clean” run, only the “clean” label $\alpha_{j^{\prime}}$. In practice, other subspaces of $\mathbf{x}^{j}_{i,l}$ may be used by the model in alternate pathways to compute the answer. To avoid this affecting the intervention, we average out all subspaces not in the intervened subspace. Letting $\overline{{\bf x}_{i,l}}$ be the average of ${\bf x}^{j}_{i,l}$ across all prompts indexed by $j$ and ${\mathbf{P}}^{+}$ be the pseudoinverse of ${\mathbf{P}}$, we intervene via the formula
@@
5.2 Decomposing Hidden States
To isolate the rough circuit for Weekdays and Months, we perform layer-wise activation patching on $40$ random pairs of prompts. The results, displayed in Appendix I, show that the circuit to compute $\gamma$ consists of MLPs on top of the $\alpha$ and $\beta$ tokens, a copy to the token before $\gamma$, and further MLPs there (roughly similar to prior work studying arithmetic circuits [43]). Moreover, fine-grained patching in Appendix I shows that there are just a few responsible attention heads for the writes to the token before $\gamma$. However, patching alone cannot tell use how or where $\gamma$ is represented.
We now introduce a new technique for empirically explaining hidden representations in algorithmic problems: explanation via regression (EVR). Given a set of tokens $T$ with a corresponding set of hidden states $X_{i,l}$, we explain the variance in $X_{i,l}$ by adding together hand-chosen functions of $t$. This gives us an explanation of what the transformation $T\rightarrow X_{i,l}$ computes. For a given choice of explanation functions $\{{\mathbf{g}}_{i}(t)\}$, the $r^{2}$ value of a linear regression from $\{{\mathbf{g}}_{i}(t)\}$ to $X_{i,l}$ gives a measure of the explained variance in $X_{i,l}$. But what functions ${\mathbf{g}}_{i}$ should we choose?
In  Fig. 8, we perform EVR on the layer 17-29 hidden states of Mistral 7B on the Weekdays task; additional deconstructions are in Appendix I. We find that a circle in $\gamma$ develops and grows in explanatory power; we plot the layer $25$ residuals after explaining with one hot functions in $\alpha$ and $\beta$ (i.e. ${\mathbf{g}}_{1}=[\alpha=0],{\mathbf{g}}_{2}=[\beta=1],{\mathbf{g}}_{3}=[\alpha=1],\ldots$) in Fig. 7 to show this incredibly clear circle in $\gamma$. This suggests that the models may be generating $\gamma$ by using a trigonometry based algorithm like the “clock” [35] or “pizza” [49] algorithm in late MLP layers.
@@ 6 Discussion
Our work proposes a significant refinement to the simple one-dimensional linear representation hypothesis. While previous work has convincingly shown the existence of one-dimensional features, we find evidence for multi-dimensional representations that are non-separable and irreducible, requiring us to generalize the notion of a feature to higher dimensions. Fortunately, we find that existing feature extraction methodologies like sparse autoencoders can readily be applied to discover multi-dimensional representations. Although multi-dimensional representations may be more complicated, we believe that uncovering the true (perhaps multi-dimensional) nature of model representations is necessary for discovering the underlying algorithms that use these representations. Ultimately, we aim to turn complex circuits in future more-capable models into formally verifiable programs [46, 10], which requires the ground truth “variables” of language models; we believe this work takes an important step towards discovering these variables. Finally, we do not anticipate adverse impacts of our work, as it focuses only on deepening our understanding of language model representations.
@@
It is unclear why we did not find more interpretable multi-dimensional features: are there truly not that many, or is our clustering technique failing to find them? Our definition for an irreducible feature (Definition 2) also had to be relaxed to hold in practice (Definition 3). Thus, although this work provides preliminary evidence for the multi-dimensional superposition hypothesis (2), it is still unclear if this theory provides the best description for the representations models use. We also did not find a small subset of MLP neurons implementing the “clock” algorithm, leaving as an open question to what extent models use multi-dimensional representations for algorithmic tasks. Finally, we only ran experiments on models up to size 8B; however, recent work [20] implies representations may become universal with growing model size.
Acknowledgments and Disclosure of Funding
@@
Appendix A Proofs
@@
We will first bound the $L_{2}$ norm of $\mathbf{z}$. If $\sigma_{n}$ is the minimum singular value of ${\mathbf{A}}$, then we have via standard singular value inequalities [18]
@@
Thus we now lower bound $\sigma_{n}$. The singular values are the square roots of the eigenvalues of the matrix ${\mathbf{A}}^{T}{\mathbf{A}}$, so we now examine ${\mathbf{A}}^{T}{\mathbf{A}}$. Since all elements of ${\mathbf{A}}$ are unit vectors, the diagonal of ${\mathbf{A}}^{T}{\mathbf{A}}$ is all ones. The off diagonal elements are dot products of pairs of $\delta$-orthogonal vectors, and so are within the range $[-\delta,\delta]$. Then by the Gershgorin circle theorem [13], all eigenvalues $\lambda_{i}$ of ${\mathbf{A}}^{T}{\mathbf{A}}$ are in the range
@@
By the JL lemma [23, 19], for any $d$ and $\delta$, we can choose $e^{Cd\delta^{2}}$ $\delta$-orthogonal unit vectors in $\mathbb{R}^{d}$ indexed as $\mathbf{v}_{i}$, for some constant $C$. Let ${\mathbf{A}}_{i}=[\mathbf{v}_{d_{\max}*i},\ldots,\mathbf{v}_{d_{\max}*i+n_{i}-1}]$ where each element in the brackets is a column. Then by construction all ${\mathbf{A}}_{i}$ are matrices composed of unique $\delta$-orthogonal vectors and there are $\frac{1}{d_{\max}}e^{Cd\delta^{2}}$ matrices ${\mathbf{A}}_{i}$.
@@
Thus ${\mathbf{A}}_{i}$ and ${\mathbf{A}}_{j}$ are $f(\delta)$-orthogonal for $f(\delta)=\delta d_{\max}/(1-\delta d_{\max})$, and so it is possible to choose $\frac{1}{d_{\max}}e^{Cd\delta^{2}}$ pairwise $f(\delta)$-orthogonal projection matrices. Remapping the variable $\delta$ with $\delta\mapsto f^{-1}(\delta)=\delta/(d_{\max}(1+\delta))$, we find that it is possible to choose $\frac{1}{d_{\max}}e^{Cd\delta^{2}/((1+\delta)^{2}d_{\max}^{2})}$ pairwise $\delta$-orthogonal projection matrices. Because $1+\delta$ is at most $2$ with $\delta\in(0,1)$, we can further simplify the exponent and find that it is possible to choose $\frac{1}{d_{\max}}e^{C(d/d_{\max}^{2})\delta^{2}/4}$ pairwise $\delta$-orthogonal projection matrices. Absorbing the $4$ into the constant $C$ finishes the proof. ∎

### v2

@@ 4 Sparse Autoencoders Find Multi-Dimensional Features
Sparse autoencoders (SAEs) deconstruct model hidden states into sparse vector sums from an over-complete basis (Bricken et al., 2023; Cunningham et al., 2023). For hidden states $X_{i,l}$, a one-layer SAE of size $m$ with sparsity penalty $\lambda$ minimizes the following dictionary learning loss (Bricken et al., 2023; Cunningham et al., 2023):
@@
We now argue that SAEs can discover irreducible multi-dimensional features by clustering ${\mathbf{D}}$. We will consider a simple form of clustering: build a complete graph on ${\mathbf{D}}$ with edge weights equal to the cosine similarity between dictionary elements, prune all edges below a threshold $T$, and then set the clusters equal to the connected components of the graph. If we now consider the spaces spanned by each cluster, they will be approximately $T$-orthogonal by construction, since their basis vectors are all $T$-orthogonal. Now, consider some irreducible two-dimensional feature ${\mathbf{f}}$; we claim that one of the clusters is likely to be exactly equal to ${\mathbf{f}}$. If ${\mathbf{D}}$ includes just two dictionary elements spanning ${\mathbf{f}}$, then these elements both must have nonzero activations post-ReLU to reconstruct ${\mathbf{f}}$ (otherwise ${\mathbf{f}}$ is a mixture). Because of the sparsity penalty in Eq. 4, this two-vector solution to reconstruct ${\mathbf{f}}$ is disincentivized, so instead the dictionary is likely to learn many elements that span ${\mathbf{f}}$. These dictionary elements will then have a high cosine similarity, and so the edges between them will not be pruned away during the clustering process; hence, they will be in a cluster.
Thus, we have a way to operationalize 2: clustering $D$ finds $T$-orthogonal subspaces, and if irreducible multi-dimensional features exist, they are likely to be equal to some of these subspaces. This suggests a natural approach to using sparse autoencoders to search for irreducible multi-dimensional features:
1. Cluster dictionary elements by their pairwise cosine similarity. We use both the simple similarity-based pruning technique described above, as well as spectral clustering; see Appendix F for details, including comments on scalability.
@@
3. Examine the resulting reconstructed activation vectors for irreducible multi-dimensional features. This step can be done manually by visually inspecting the PCA projections for known irreducible multi-dimensional structures (e.g. circles, see Fig. 10) or automatically by passing the PCA projections to the tests for Definition 3.
Figure 3: Mixture index and separability index of GPT-2 features. Features from Fig. 1, which we had manually identified, score highly as candidate multidimensional features with these metrics.
@@
This method succeeds on toy datasets of synthetic irreducible multi-dimensional features; see Appendix D. [^3] Experiment code: https://github.com/JoshEngels/MultiDimensionalFeatures We apply this method to language models using GPT-2 (Radford et al., 2019) SAEs trained by Bloom (2024) for every layer and Mistral 7B (Jiang et al., 2023) SAEs that we train on layers 8, 16, and 24 (training details in Appendix E).
@@
For each cluster of GPT-2 SAE features, we take the reconstructed activations and project them onto PCA components 1-2, 2-3, 3-4, and 4-5 (or fewer if there are fewer features in the cluster) and measure the separability index and $\epsilon$-mixture index of each 2D point cloud as described in Section B.2. The mean scores across these planes are a computationally tractable approximation of Definition 3. We plot these mean scores in Fig. 3, and find that the features which we had manually identified in Fig. 1 are among the top scoring features along both measures of irreducibility. Thus, our theoretical tests can indeed be used to find interpretable irreducible features.
@@ 5 Circular Representations in Large Language Models
In this section, we seek tasks in which models use the multi-dimensional features we discovered in Section 4, thereby providing evidence that these representations are indeed the fundamental unit of computation for some problems. Inspired by prior work studying circular representations in modular arithmetic (Liu et al., 2022), we define two prompts that represent “natural” modular arithmetic tasks:
@@
For Weekdays, we range over the $7$ days of the week and durations between $1$ and $7$ days to get $49$ prompts. For Months, we range over the $12$ months of the year and durations between $1$ and $12$ months to get $144$ prompts. Mistral 7B and Llama 3 8B (AI@Meta, 2024) achieve reasonable performance on the Weekdays task and excellent performance on the Months task (measured by comparing the highest logit valid token against the ground truth answer), as summarized in Table 1. Interestingly, although these problems are equivalent to modular arithmetic problems $\alpha+\beta\equiv\;?\ (\mathrm{mod}\ m)$ for $m=7,12$, both models get trivial accuracy on plain modular addition prompts, e.g. “5+3(mod $5+3\ (\mathrm{mod}\ 7)\equiv$”. Finally, although GPT-2 has circular representations, it gets trivial accuracy on Weekdays and Months.
@@ 5.1 Intervening on Circular Day and Month Representations
We now experiment with intervening on these circular representations. We base our experiments on the common interpretability technique of activation patching, which replaces activations from a “dirty” run of the model with the corresponding activations from a “clean” run (Zhang & Nanda, 2023). Activation patching empirically tests whether a specific model component, position, and/or representation has a causal influence on the model’s output. We employ a custom subspace patching method to allow testing for whether a specific circular subspace of a hidden state is sufficient to causally explain model output. Specifically, our patching technique relies on the following steps (visualized in Fig. 6):
@@
2. Intervene on the subspace: Say our initial prompt had $\alpha=\alpha_{j}$ and we are intervening with $\alpha=\alpha_{j^{\prime}}$. In this step, we replace the model’s projection on the subspace ${\mathbf{P}}\cdot{\mathbf{W}}_{i,l}$, which will be close to $\texttt{circle}(\alpha_{j}$), with the “clean” point $\texttt{circle}(\alpha_{j^{\prime}})$. Note that we do not use the hidden state $\mathbf{x}^{j^{\prime}}_{i,l}$ from the “clean” run, only the “clean” label $\alpha_{j^{\prime}}$. In practice, other subspaces of $\mathbf{x}^{j}_{i,l}$ are used by the model in alternate pathways to compute the answer, so if we just intervene on the circular subspace the logit difference effect is not as pronounced. To solve this problem, we average out the portion of the activation not in the intervened subspace. Letting $\overline{{\bf x}_{i,l}}$ be the average of ${\bf x}^{j}_{i,l}$ across all prompts indexed by $j$ and ${\mathbf{P}}^{+}$ be the pseudoinverse of ${\mathbf{P}}$, we intervene via the formula
@@
5.2 Uncovering Output Representations Using Regression
So far, we have focused on examining and intervening on the representation for $\alpha$, which was conveniently present as a circle in the top PCA components on top of the $\alpha$ token. In this section, we will examine how the generated output, $\gamma$, is represented.
First, to isolate the rough circuit for Weekdays and Months, we perform layer-wise activation patching on $40$ random pairs of prompts. The results, displayed in Appendix J, show that the circuit to compute $\gamma$ consists of MLPs on top of the $\alpha$ and $\beta$ tokens, a copy to the token before $\gamma$, and further MLPs there (roughly similar to what Stolfo et al. (2023) find in prior work studying arithmetic circuits). Thus, we know where to look for a representation of $\gamma$: in the second half of the layers on the token before $\gamma$. However, patching alone cannot tell use how $\gamma$ is represented.
@@ 5.2 Uncovering Output Representations Using Regression
Unlike $\alpha$, $\gamma$ has no obvious circular (or linear) pattern in the top PCA components on these layers. To determine the representation for $\gamma$, we introduce a more powerful technique we call Explanation via Regression (EVR): given a set of token sequences with a corresponding set of hidden states $X_{i,l}$, we choose a set of interpretable explanation functions of the input tokens $\{{\mathbf{g}}_{j}(t)\}$. The $r^{2}$ value of a linear regression from $\{{\mathbf{g}}_{j}(t)\}$ to $X_{i,l}$ tells us how much of the variance in the activations the $\{{\mathbf{g}}_{j}(t)\}$ explain, and conversely the residuals show the exact components of the representation we have yet to explain.
More details and experiments using EVR to completely break down language model states are in Appendix J. Here, we use EVR to determine the representation for $\gamma$ by plotting the top two PCA components of the layer $25$ Mistral 7B activations after subtracting the components that can be explained using a regression with one hot functions in $\alpha$ and $\beta$ (i.e. ${\mathbf{g}}_{1}=[\alpha=0],{\mathbf{g}}_{2}=[\beta=1],{\mathbf{g}}_{3}=[\alpha=1],\ldots$). The result, shown in Fig. 8, is an incredibly clear circle in $\gamma$, which suggests that the model’s generated representation of $\gamma$ lies along a circle. A simple PCA projection was not enough to find this result because the representation for $\gamma$ has interference from $\alpha$ and $\beta$, which the EVR removes. This suggests that the models may be generating $\gamma$ by using a trigonometry based algorithm like the “clock” Nanda et al. (2023a) or “pizza” Zhong et al. (2024) algorithm in late MLP layers.
5.3 Continuity of Circular Representations
Figure 9: Layer $30$ Mistral 7B activations for [very early/very late] on [Monday/Tuesday/…/Sunday], plotted projected into the PCA plane for [Monday/Tuesday/…/Sunday].
In past sections, the representations of the interpretable numeric quantities we have discovered have been mostly discontinuous; that is, the days of the week and months of the year in Fig. 1 and Fig. 15 are clustered at the vertices of a heptagon and dodecagon, and there is nothing "between" adjacent weekdays or months along the circle. Although continuity of the representation is not a requirement of Definition 3, it would further decrease the $\epsilon$-mixture index, and would also increase our subjective perception of the circular feature as an intrinsic model feature representing a continuous quantity (time). Thus, we create a synthetic dataset containing the text "[very early/very late] on [Monday/Tuesday/…/Sunday]" and simply plot the projections of the layer $30$ activations into the top two PCA components of the activations of [Monday/Tuesday/…/Sunday]. The results, shown in Fig. 9, show that Mistral 7B indeed can map intermediate quantities to their expected place in the circle: the very early and very late version of each weekday are more towards the last and the next weekday along the circle, respectively. We show similar results for "[morning/evening] on [Monday/Tuesday/…/Sunday]" in Appendix Fig. 19.
@@ 6 Discussion
Our work proposes a significant refinement to the simple one-dimensional linear representation hypothesis. While previous work has convincingly shown the existence of one-dimensional features, we find evidence for irreducible multi-dimensional representations, requiring us to generalize the notion of a feature to higher dimensions. Fortunately, we find that existing unsupervised feature extraction methodologies like sparse autoencoders can readily be applied to discover multi-dimensional representations. However, we think our work raises interesting questions about whether individual SAE features are appropriate “mediators” (Mueller et al., 2024) for understanding model computation, if some features are in fact multi-dimensional. Although taking a multi-dimensional representation perspective may be more complicated, we believe that uncovering the true (perhaps multi-dimensional) nature of model representations is necessary for discovering the underlying algorithms that use these representations. Ultimately, our field aims to turn complex circuits in future more-capable models into formally verifiable programs (Tegmark & Omohundro, 2023; Dalrymple et al., 2024), which requires the ground truth “variables” of language models; we believe this work takes an important step towards discovering these variables.
@@
It is unclear why we did not find more interpretable multi-dimensional features. We are unsure if we are failing to interpret some of the high-scoring multi-dimensional features, if our clustering technique is not powerful enough to find some features, or if there are truly not that many. Our definition for an irreducible feature (Definition 2) also had to be relaxed to hold in practice (Definition 3). Thus, although this work provides preliminary evidence for the multi-dimensional superposition hypothesis (2), it is still unclear if this theory provides the best description for the representations models use. We also did not find a small subset of MLP neurons implementing the “clock” algorithm, leaving as an open question to what extent models use multi-dimensional representations for algorithmic tasks. Finally, we only ran experiments on models up to size 8B; however, recent work (Huh et al., 2024) implies representations may become universal with growing model size.
Acknowledgments
@@
Appendix A Multi-Dimensional Feature Capacity
The Johnson-Lindenstrauss (JL) Lemma (Johnson & Lindenstrauss, 1984) implies that we can choose $e^{Cd\delta^{2}}$ pairwise one-dimensional $\delta$-orthogonal vectors to satisfy 1 for some constant $C$, thus allowing us to build the model’s feature space with a number of one-dimensional $\delta$-orthogonal features exponential in $d$. We now prove a similar result for low-dimensional projections (the main idea of the proof is to combine $\delta$-orthogonal vectors as guaranteed from the JL lemma):
Theorem 1.
For any $d^{\prime}$ and $\delta$, it is possible to choose $\frac{1}{d_{\max}}e^{C_{1}(d/d^{\prime 2})\delta^{2}}$ pairwise $\delta$-orthogonal matrices ${\mathbf{A}}_{i}\in\mathbb{R}^{n_{i}\times d^{\prime}}$ for some constant $C_{1}$. Furthermore, it is not possible to choose more than $e^{C_{2}(d-d_{\max}\delta\log(\frac{1}{\delta}))}$ for some constant $C_{2}$.
@@ Appendix A Multi-Dimensional Feature Capacity
We will first bound the $L_{2}$ norm of $\mathbf{z}$. If $\sigma_{n}$ is the minimum singular value of ${\mathbf{A}}$, then we have via standard singular value inequalities (Higham, 2021)
@@
Thus we now lower bound $\sigma_{n}$. The singular values are the square roots of the eigenvalues of the matrix ${\mathbf{A}}^{T}{\mathbf{A}}$, so we now examine ${\mathbf{A}}^{T}{\mathbf{A}}$. Since all elements of ${\mathbf{A}}$ are unit vectors, the diagonal of ${\mathbf{A}}^{T}{\mathbf{A}}$ is all ones. The off diagonal elements are dot products of pairs of $\delta$-orthogonal vectors, and so are within the range $[-\delta,\delta]$. Then by the Gershgorin circle theorem (Gershgorin, 1931), all eigenvalues $\lambda_{i}$ of ${\mathbf{A}}^{T}{\mathbf{A}}$ are in the range
@@
By the JL lemma (Johnson & Lindenstrauss, 1984; (), https://mathoverflow.net/users/2554/bill johnson), for any $d$ and $\delta$, we can choose $e^{Cd\delta^{2}}$ $\delta$-orthogonal unit vectors in $\mathbb{R}^{d}$ indexed as $\mathbf{v}_{i}$, for some constant $C$. Let ${\mathbf{A}}_{i}=[\mathbf{v}_{d_{\max}*i},\ldots,\mathbf{v}_{d_{\max}*i+n_{i}-1}]$ where each element in the brackets is a column. Then by construction all ${\mathbf{A}}_{i}$ are matrices composed of unique $\delta$-orthogonal vectors and there are $\frac{1}{d_{\max}}e^{Cd\delta^{2}}$ matrices ${\mathbf{A}}_{i}$.
@@
Thus ${\mathbf{A}}_{i}$ and ${\mathbf{A}}_{j}$ are $f(\delta)$-orthogonal for $f(\delta)=\delta d_{\max}/(1-\delta d_{\max})$, and so it is possible to choose $\frac{1}{d_{\max}}e^{Cd\delta^{2}}$ pairwise $f(\delta)$-orthogonal projection matrices. Remapping the variable $\delta$ with $\delta\mapsto f^{-1}(\delta)=\delta/(d_{\max}(1+\delta))$, we find that it is possible to choose $\frac{1}{d_{\max}}e^{Cd\delta^{2}/((1+\delta)^{2}d_{\max}^{2})}$ pairwise $\delta$-orthogonal projection matrices. Because $1+\delta$ is at most $2$ with $\delta\in(0,1)$, we can further simplify the exponent and find that it is possible to choose $\frac{1}{d_{\max}}e^{C(d/d_{\max}^{2})\delta^{2}/4}$ pairwise $\delta$-orthogonal projection matrices. Absorbing the $4$ into the constant $C$ finishes the proof of the lower bound.
For the upper bound, we can proceed much more simply. Consider $k$ pairwise $\delta$-orthogonal matrices $A_{i}\in\mathbb{R}^{d^{\prime}}$. Since these matrices are full rank, their column spaces each parameterize a subspace of dimension $d^{\prime}$, and so by a result from (Alon, 2003) it is possible to choose $e^{Cd\textquoteright\delta^{2}\log(\frac{1}{\delta})}$ almost orthogonal vectors in this subspace. Furthermore, by our definition of $\delta$-orthogonal matrices, all pairs of these vectors between subspaces will be $\delta$-orthogonal. Finally, again by (Alon, 2003) we cannot have more than $e^{Cd\delta^{2}\log(\frac{1}{\delta})}$ $\delta$-orthogonal vectors overall, so we have that
$ke^{Cd_{\max}\delta^{2}\log(\frac{1}{\delta})}$ $<e^{Cd\delta^{2}\log(\frac{1}{\delta})}$
and simplfying,
$k$ $<e^{C(d-d_{\max})\delta^{2}\log(\frac{1}{\delta})}$
These results imply that models can still represent an exponential number of higher dimensional features. However, there is a large exponential gap between the lower and upper bound we have shown. If the lower bound is reasonably tight, then this would mean that models would be highly incentivized to fit features within the smallest dimensional space possible, suggesting a reason for recent work showing interesting compressed encodings of multi-dimensional features in toy problems (Morwani et al., 2023).
Note that the proof assumes the “worst case” scenario that all of the features are dimension $d_{\max}$, while in practice many of the features may be $1$ or low dimensional, so the effect on the capacity of a real model that represents multi-dimensional features is unlikely to be this extreme.
Finally, we note that the dictionary learning literature may have discovered similar results in the past (which we were unaware of), see Foucart & Rauhut (2013).